# HiLoMix: Robust High- and Low-Frequency Graph Learning Framework for Mixing Address Association

### v1

@@ CONCLUSION
In this work, we propose HiLoMix, a novel graph-based learning framework tailored for address association in mixing transactions. We model diverse account interactions in Tornado Cash as a heterogeneous graph, enabling analysis from a graph perspective, thus addressing the graph sparsity problem. To alleviate the limited supervision and label noise problem, we introduce frequency-aware graph contrastive learning alongside confidence-based label-weighted supervised learning, jointly optimizing high- and low-pass graph neural networks for robust node representations. Then a stacking framework is adopted to integrate the complementary knowledge captured by the two GNNs for the final prediction. Extensive experiments demonstrate that HiLoMix consistently outperforms existing approaches on the mixing address association task. Furthermore, we curate an up-to-date ground-truth dataset to facilitate future research on mixing address association.

### v2

@@ Conclusion
In this work, we propose HiLoMix, a novel graph-based learning framework tailored for mixing address association. By modeling diverse account interactions in Tornado Cash as HAMIG, we effectively address the issue of graph sparsity. To alleviate the limited supervision and label noise problem, we introduce frequency-aware graph contrastive learning alongside confidence-based label-weighted supervised learning, jointly optimizing high-pass and low-pass GNNs for robust node representations. Extensive experiments demonstrate that HiLoMix consistently outperforms existing approaches on the mixing address association task. Furthermore, we curate an up-to-date ground-truth dataset to support and facilitate future research in this field.
Acknowledgements
This work was supported by Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing and the National Natural Science Foundation of China (U22B2032).